Two masses 10kg and 20kg are connected by a massless string. A force of 200N acts on 20kg mass. At that instant when 10kg mass has an acceleration 12m/s^2 the energy stored in the spring (k=2400N/m) will be

Grade:12th pass

1 Answers

Khimraj
5 years ago
Acceleration of 10kg mass =12m/s²
Force=ma
F=10x12=120N will act .
Net force on 20kg mass will be :
200N-120 N=80N

a=F/m
=80/20
=4m/s²
∴ The acceleration of 20kg mass is 4m/s²